OpenStudy (anonymous):

What is the value of n in the equation 3n + 2(n + 2) = 9n + 12? −4 −2 2 4

OpenStudy (anonymous):

are those choices?

Mehek (mehek14):

have you tried solving it @Suspect ?

OpenStudy (anonymous):

So the original equation is 3n + 2(n+2) = 9n + 12 So the first step is to distribute the 2 in 2(n+2). So that would be 2n +4. The equation would then look like this : 3n + 2n + 4 = 9n +12. So then combining the variable n on the left side of the equation. That would give us 5n. The equation would then look like this : 5n + 4 = 9n + 12. The goal is to get all of the variables on one side, with the constants on the other. So, personally, I would move the 5n and the 12 to the opposite side of the equal sign. That would then make the equation look like this: -12 + 4 = 9n - 5n. The positive and negative signs change when the numbers/variables change sides on the equations, so that's why the 12 now has a negative sign in front of it. Same with the 5n. Now, we just have to combine the constants and the variables. That would make the equation look like this: -8 = 4n. Then, the goal is to get the variable alone, so that would mean dividing the 4 away from the n. When we do that, both sides of the equation must be divided by the same thing, so the -8 would be divided by the 4 also. The equation would look like this: -2 = n. So, the answer to the equation solving for n, would be -2. The second choice of the choices, which I am assuming is choice B.

OpenStudy (anonymous):

Does that help? @Suspect